### v1

# JEL classification: C11, C32, C55, E50

Abstract
We propose a high-dimensional structural vector autoregression framework capable of accommodating a large number of linear inequality restrictions on impact impulse responses, structural shocks, and their element-wise products. Combining impact- and shock-inequality restrictions can be flexibly used to sharpen inference and to disentangle structurally interpretable shocks through sign and/or shock constraints. To estimate the model, we develop a highly efficient sampling algorithm that scales well with model dimension and the number of inequality restrictions on impact responses, as well as structural shocks. It remains computationally feasible even when existing algorithms may break down. To demonstrate the practical utility of our approach, we identify five structural shocks and examine the dynamic responses of thirty macroeconomic variables, highlighting the model’s flexibility and feasibility in complex empirical settings. We provide empirical evidence that financial shocks are the most important driver of the dynamics of the business cycle.
Keywords: Structural Identification, Sign Restrictions, Narrative Restrictions, Large BVARs
JEL classification: C11, C32, C55, E50
1 Introduction
Uhlig, (2005) popularizes the use of sign restrictions on impulse response functions, offering a flexible and theory-consistent approach to identifying structural shocks in Vector Autoregression (VAR) models. Instead of relying on often implausible zero restrictions, sign restrictions constrain the signs of impulse responses based on economic intuition. Building on this framework, Antolín-Díaz and Rubio-Ramírez, (2018) introduce narrative restrictions, which incorporate external prior information about specific historical episodes-such as well-documented policy interventions or economic events-to further sharpen set identification. By conditioning structural shocks on particular time periods, narrative restrictions enable researchers to embed economically meaningful prior knowledge directly into the model, enhancing interpretability and credibility of the identified shocks. Using narrative restrictions have become popular recently. [^1] Important examples that use narrative restrictions include Zeev, (2018), Altavilla et al., (2019), Furlanetto and Robstad, (2019), Cheng and Yang, (2020), Kilian and Zhou, (2020), Kilian and Zhou, (2022), Laumer, (2020), Redl, (2020), Zhou, (2020), Antolin-Diaz et al., (2021), Caggiano et al., (2021), Larsen, (2021), Ludvigson et al., (2021), Maffei-Faccioli and Vella, (2021), Berger et al., (2022), Fanelli and Marsi, (2022), Inoue and Kilian, (2022), Badinger and Schiman, 2023a , Berthold, (2023), Caggiano and Castelnuovo, (2023), Conti et al., (2023), Harrison et al., (2023), Herwartz and Wang, (2023) Neri, (2023), Reichlin et al., (2023), Ascari et al., (2023), Boer et al., (2024) and Rüth and Van der Veken, (2023). However, in most applications researcher use rather small dimensional VARs with narrative restrictions. Furthermore, the implementation of narrative restriction rely on accept-and-reject algorithms (see, e.g., Antolín-Díaz and Rubio-Ramírez, (2018) or Giacomini et al., (2021)), which may become computationally burdensome as the number of restrictions increases. Therefore, applications with narrative restriction are limited to VARs with a small number of narrative restrictions, typically for just a single structural shock.
Since the influential work of Bańbura et al., (2010), large VARs have become a cornerstone of modern structural analysis and forecasting. By incorporating dozens-or even hundreds-of variables, large VARs offer a powerful solution to a key limitation of smaller models: omitted variable bias. This bias can lead to distorted forecasts, unreliable policy conclusions, and incorrect identification of structural shocks. Expanding the information set helps capture the economic environment more comprehensively and reduces the risk of excluding relevant dynamics. As emphasized by Hansen and Sargent, (2019) and Lippi and Reichlin, (1993, 1994), using a narrower information set than that available to economic agents can make a model non-fundamental, thus preventing accurate recovery of structural shocks. Building on this foundation, numerous studies-including Carriero et al., (2009); Giannone et al., (2015); Jarocinski and Maćkowiak, (2017); Huber and Feldkircher, (2019); Chan et al., (2024)-have explored various methodological and empirical extensions. Overall, large VARs provide a flexible framework for structural inference in high-dimensional settings to address informational deficiencies.
In this paper, we propose a high-dimensional Structural Vector Autoregression (SVAR) framework capable of accommodating a large number of linear inequality constraints on impact impulse responses, structural shocks, and their element-wise products. While previous literature has primarily focused on small narrative SVAR models, our approach extends these methods by enabling the inclusion of a much larger set of restrictions as well as a greater number of variables. The ability to incorporate a broader set of inequality restrictions enhances the structural set identification, allowing for more precise inference, see Antolín-Díaz and Rubio-Ramírez, (2018). Moreover, in cases where impact sign restrictions alone cannot easily distinguish between different structural shocks, additional sign restrictions directly on the structural shocks provide a flexible way to separate them.
Recently, the assumption that VAR errors follow a factor structure has gained popularity, with latent factors being interpreted as structural shocks (see, for instance, Korobilis, (2022), Banbura et al., (2023), Chan et al., (2022),Hauzenberger et al., (2022), Gambetti et al., (2023), Pfarrhofer and Stelzer, (2024), Prüser, (2024), Korobilis, (2025) and Chan and Qi, (2025)). We build on this literature, leveraging the factor structure for several key advantages. First, from a computational perspective, the factor structure enables equation-by-equation estimation, simplifying the estimation process and allowing for the inclusion of a rich set of variables in the model. Second, as the number of variables increases (including different measures of prices and output), it is reasonable to assume that the number of structural shocks is smaller than the number of variables, making a factor-based approach well-suited for high-dimensional settings. Third, and particularly important for our paper, the factors have a clear interpretation as structural shocks. This provides a natural framework for incorporating narrative restrictions in the form of prior distributions on the structural shocks, which allows to develop a simple yet efficient algorithm.
Our approach deviates from existing methods such as those in Antolín-Díaz and Rubio-Ramírez, (2018) and Giacomini et al., (2021), which typically incorporate narrative restrictions through a pseudo-likelihood. Instead, imposing shock sign restrictions via prior distributions allows us to introduce a new sampling algorithm that offers significant computational efficiency by ensuring that each draw is accepted with certainty. This feature enables us to handle a large number of restrictions on both impact responses and structural shocks, overcoming computational challenges often encountered with accept-and-reject algorithms. Our approach therefore represents a major advancement in the scalability and feasibility of SVAR models, particularly in high-dimensional settings where existing algorithms may be very slow or even infeasible.
In an empirical application, we demonstrate the practical utility of our approach by identifying five structural shocks in a large VAR. In this respect, we resemble Stock and Watson, (2012) and Hou, (2024) and estimate 30 variable VAR and disentangle structurally interpretable shocks by means of impact and shock sign restrictions. Importantly, we show that shock sign restrictions can help to separate structural shocks and to overall sharpen identification. In our empirical illustration, we provide evidence suggesting that financial risk shocks are a key driver of business cycle dynamics. This example underscores the flexibility and scalability of our methodology, which is capable of handling high-dimensional data and a large number of inequality restrictions on the impact responses as well as on the structural shocks.
The remainder of this paper is organized as follows. Section 2 lays out and discusses the econometric framework. Section 3 applies the model to the study of Stock and Watson, (2012) to identify and investigate the effects of five structural shocks on the economy. Section 4 concludes.
2 A large structural VAR with factor structure
Let $\bm{y}_{t}=(y_{1,t},\dots,y_{n,t})^{\prime}$ represent an $n\times 1$ vector of endogenous variables at time $t$. The model can be expressed as:
$\bm{y}_{t}=\bm{b}_{0}+\bm{B}_{1}\bm{y}_{t-1}+\cdots+\bm{B}_{p}\bm{y}_{t-p}+\bm{u}_{t},$ (1)
where the error term $\bm{u}_{t}$ is decomposed as
$\bm{u}_{t}=\bm{L}\bm{f}_{t}+\bm{v}_{t},$ (2)
with $\bm{v}_{t}\sim N(\bm{0},\bm{\Sigma})$, where $\bm{\Sigma}=\text{diag}(\sigma_{1}^{2},\dots,\sigma_{n}^{2})$, and $\bm{f}_{t}$ is an $r\times 1$ vector of factors such that $\bm{f}_{t}\sim N(\bm{0},\bm{I})$. Concisely, the system can be reformulated as:
$\bm{y}_{t}=(\bm{I}_{n}\otimes\bm{x}^{\prime}_{t})\bm{\beta}+\bm{L}\bm{f}_{t}+\bm{v}_{t},$ (3)
where $\bm{I}_{n}$ is the identity matrix of dimension $n$, $\otimes$ represents the Kronecker product, and $\bm{\beta}=\text{vec}([\bm{b}_{0},\bm{B}_{1},\dots,\bm{B}_{p}]^{\prime})$. The vector $\bm{x}_{t}=(1,\bm{y}^{\prime}_{t-1},\dots,\bm{y}^{\prime}_{t-p})^{\prime}$ of dimension $k=1+np$ contains an intercept and lagged values. The idiosyncratic component $\bm{v}_{t}$ accounts for measurement error or other idiosyncratic noise. In contrast the $r$ latent factors, $\bm{f}_{t}$ influencing multiple variables in the system and hence have the interpretation as structural shocks. More formally we multiply the model with the generalized inverse of $\bm{L}$ to obtain:
$\bm{A}\bm{y}_{t}=\bm{B}\bm{x}_{t}+\bm{f}_{t}+\bm{A}\bm{v}_{t},$ (4)
where $\bm{A}=(\bm{L}^{\prime}\bm{L})^{-1}\bm{L}^{\prime}$ and $\bm{B}=(\bm{A}\bm{b}_{0},\bm{A}\bm{B}_{1},\dots,\bm{A}\bm{B}_{p})$. Given that $\bm{v}_{t}$ is uncorrelated noise, the Central Limit Theorem (see Bai, (2003)) implies that $\bm{A}\bm{v}_{t}\to 0$ as $n\to\infty$. Hence, we can write
$\bm{f}_{t}\approx\bm{A}\bm{y}_{t}-\bm{B}\bm{x}_{t}.$ (5)
Thus, $\bm{v}_{t}$ is treated as noise shocks without structural meaning, while $\bm{f}_{t}$ provides a projection of structural shocks in $\mathbb{R}^{r}$. Consequently, this formulation supports structural analysis using standard methods, including impulse response functions (e.g., Forni et al., (2019), Korobilis, (2022), Chan et al., (2022)).
Under the assumption of uncorrelated $\bm{f}_{t}$ and $\bm{v}_{t}$, the conditional variance of $\bm{u}_{t}$ is given by:
$\text{Var}(\bm{u}_{t}|\bm{\Sigma},\bm{L})=\bm{L}\bm{L}^{\prime}+\bm{\Sigma}.$ (6)
To separately identify the common and idiosyncratic components, we adopt the condition $r\leq(n-1)/2$ from Anderson and Rubin, (1956). Under this restriction, for any observationally equivalent model $(\bm{L},\bm{\Sigma})$, it holds that $\bm{L}\bm{L}^{\prime}=\bm{L}^{*}\bm{L}^{*^{\prime}}$ and $\bm{\Sigma}=\bm{\Sigma}^{*}$.
However, without additional restrictions, $\bm{L}$ is not identified. Any orthogonal matrix $\bm{Q}\in\mathcal{O}(r)$, where $\mathcal{O}(r)=\{\bm{Q}\in\mathbb{R}^{r\times r}:\bm{Q}\bm{Q}^{\prime}=\bm{I}_{m}\}$, generates an equivalent model via the transformation $\tilde{\bm{L}}\tilde{\bm{f}}_{t}=\bm{L}\bm{Q}\bm{Q}^{\prime}\bm{f}_{t}$. In this paper we achive set identification by placing inequality restrictions on $\L$, $\bm{f}$ as well as on the product of both.
2.1 Linear inequality restrictions via prior distributions
We consider inequality restriction on the $\bm{L}$, $\bm{f}_{t}$, and on their product. [^2] In many cases we have a strong consensus in economic theory about the signs of impulse responses at impact but not at longer horizons (see, e.g., Canova and Paustian, (2011)). We can express these inequality restrictions as follows:
$\ell_{i}^{L}$ $<\bm{R}_{i}^{L}\bm{l}_{i}^{\prime}<\upsilon_{i}^{L},$ (7)
$\ell_{\tilde{t}}^{f}$ $<\bm{R}_{\tilde{t}}^{f}\bm{f}_{\tilde{t}}<\upsilon_{\tilde{t}}^{f},$ (8)
$\ell_{i\tilde{t}}^{Lf}$ $<\bm{R}_{i\tilde{t}}^{Lf}(\bm{l}_{i}^{\prime}\odot\bm{f}_{\tilde{t}})<\upsilon_{i\tilde{t}}^{Lf},$ (9)
where $\bm{l}_{i}$ denote the elements of $\bm{L}$ in the $i-$th equation, $\bm{R}_{i}^{L}$ is $r_{i}^{L}\times r$, $\bm{R}_{\tilde{t}}^{f}$ is $r_{\tilde{t}}^{f}\times r$, $\bm{R}_{i\tilde{t}}^{Lf}$ is $r_{i\tilde{t}}^{Lf}\times r$ and $\tilde{t}\in t=1,\dots,T$. We assume that $\ell_{i}^{L}<\upsilon_{i}^{L}$, $\ell_{\tilde{t}}^{f}<\upsilon_{\tilde{t}}^{f}$ and $\ell_{i\tilde{t}}^{Lf}<\upsilon_{i\tilde{t}}^{Lf}$. Furthermore, we allow elements in the lower and upper bound to be $\infty$- ∞ or $\infty$ to indicate that some inequity restrictions are single-sided. In many applications $\bm{R}_{i}^{L}$, $\bm{R}_{\tilde{t}}^{f}$ and $\bm{R}_{i\tilde{t}}^{Lf}$ will be equal to $\bm{I}_{r}$. The first set of inequality restrictions can be used to implement sign restrictions on $\bm{L}$ as in Korobilis, (2022). The second set of inequality restriction can be used to implement sign restrictions on the structural shocks $\bm{f}_{t}$ at specific time periods as in Antolín-Díaz and Rubio-Ramírez, (2018). Finally, the last set of inequality restrictions can be used to implement a magnitude restrictions on the product of both factors loadings and structural shocks as in Badinger and Schiman, 2023b . In particular, Badinger and Schiman, 2023b impose that one particular shock explains more then half of the unexplained movement of a certain variable at a specific period. This provides a flexible framework for the identification of structural shocks. Instead of only relaying on impact sign restrictions, researchers can also use shock sign restrictions to distinguish between different structural shocks. Or even use a single magnitude restriction as in Badinger and Schiman, 2023b to separate one shock from the others.
We implement our inequality restrictions using truncated Gaussian prior distributions. In particular we assume
$\bm{l}_{i}$ $\sim N(\bm{l}_{0,i},\bm{V}_{\bm{l}_{i}})\bm{1}(\ell_{i}^{L}<\bm{R}_{i}^{L}\bm{l}_{i}^{\prime}<\upsilon_{i}^{L})\bigcup_{\tilde{t}\in\tilde{T}}\bm{1}(\ell_{i\tilde{t}}^{Lf}<\bm{R}_{i\tilde{t}}^{Lf}(\bm{l}_{i}^{\prime}\odot\bm{f}_{\tilde{t}})<\upsilon_{i\tilde{t}}^{Lf}),$ (10)
$\bm{f}_{t}$ $\sim N(\bm{0},\bm{I})\bm{1}(\ell_{t}^{f}<\bm{R}_{t}^{f}\bm{f}_{t}<\upsilon_{t}^{f})\bm{1}(\ell_{it}^{Lf}<\bm{R}_{it}^{Lf}(\bm{l}_{i}^{\prime}\odot\bm{f}_{t})<\upsilon_{it}^{Lf},$ (11)
where $\tilde{T}$ denotes the set of periods in which a restrictions is specified. We complete our model specification by describing the further prior distributions. For the variance terms of the noise we assume $\sigma_{j}^{2}\sim\mathcal{IG}(\alpha_{0},\beta_{0})$. In our empirical application, we set $\bm{l}_{0,i}=\bm{0}$, $\bm{V}_{\bm{l}_{i}}=10\times\bm{I}_{r}$ and $\alpha_{0}=\beta_{0}=0$. In high-dimensional settings such as large VARs, it is important to impose shrinkage prior to avoid overfitting. We follow Korobilis, (2022) and use the horseshoe prior proposed by Carvalho et al., (2010). Let $\bm{\beta}_{i}$ the VAR coefficients in the $i-$th equation, $i=1,\dots,n$ and $\beta_{i,j}$, the $j-$th coefficient in the $i-$th equation. Consider the prior for $\beta_{i,j}$, $i=1,\dots,n$ and $j=2,\dots,k$:
$\beta_{i,j}|\lambda_{i},\psi_{i,j}$ $\sim N(0,\lambda_{i}\psi_{i,j}),$ (12)
$\sqrt{\psi_{i,j}}$ $\sim C^{+}(0,1),$ (13)
$\sqrt{\lambda_{i}}$ $\sim C^{+}(0,1),$ (14)
where $C^{+}(0,1)$ denotes the standard half-Cauchy distribution. The hyperparameter $\lambda_{i}$ is the global variance components that are common to all elements in $\bm{\beta}_{i}$ , whereas each $\psi_{i,j}$ is a local variance component specific to the coefficients $\beta_{i,j}$. To facilitate sampling, we follow Makalic and Schmidt, (2015) and use the following latent variables representations of the half-Cauchy distributions:
$(\psi_{i,j}|z_{\psi_{i,j}})$ $\sim\mathcal{IG}(1/2,1/z_{\psi_{i,j}}),\quad z_{\psi_{i,j}}\sim\mathcal{IG}(1/2,1),$ (15)
$(\lambda_{i}|z_{\lambda_{i}})$ $\sim\mathcal{IG}(1/2,1/z_{\lambda_{i}}),\quad z_{\lambda_{i}}\sim\mathcal{IG}(1/2,1),$ (16)
for $i=1,\dots,n$ and $j=2,\dots,k$.
2.2 Gibbs Sampler
In this section, we develop an efficient posterior sampler to estimate the model. To impose the inequality restrictions on the factors or factor loadings we use the sampler from Botev, (2017). In contrast to existing rejecting sampling approaches (see, e.g. Antolín-Díaz and Rubio-Ramírez, (2018) or Giacomini et al., (2021) ) each draw is accepted with certainty. Posterior draws can be obtained by sampling sequentially from the conditional distributions:
  1. 1.
    $p(\bm{f}|\bm{y},\bm{\beta},\bm{L},\bm{\Sigma},\bm{\lambda},\bm{\psi},\bm{z}_{\lambda},\bm{z}_{\bm{\psi}})=p(\bm{f}|\bm{y},\bm{\beta},\bm{L},\bm{W},\bm{\Sigma})$;
  2. 2.
    $p(\bm{L}|\bm{y}\bm{\beta},,\bm{f},\bm{\Sigma},\bm{\lambda},\bm{\psi},\bm{z}_{\lambda},\bm{z}_{\bm{\psi}})=\prod_{i=1}^{n}=p(\bm{l}_{i}|\bm{y}_{i},\bm{\beta}_{i},\bm{f},\bm{\sigma}^{2}_{i})$
  3. 3.
    $p(\bm{\beta}|\bm{y},\bm{L},\bm{f},\bm{\Sigma},\bm{\lambda},\bm{\psi},\bm{z}_{\lambda},\bm{z}_{\bm{\psi}})=\prod_{i=1}^{n}=p(\bm{\beta}_{i}|\bm{y}_{i},\bm{l}_{i},\bm{f},\bm{\sigma}^{2}_{i})$
  4. 4.
    $p(\bm{\Sigma}|\bm{y},\bm{\beta},\bm{L},\bm{f},\bm{\lambda},\bm{\psi},\bm{z}_{\lambda},\bm{z}_{\bm{\psi}})=\prod_{i}^{n}p(\sigma_{i}^{2}|\bm{y}_{i},\bm{f}_{i},\bm{l}_{i},\bm{\beta}_{i})$;
  5. 5.
    $p(\bm{\lambda}|\bm{y},\bm{\beta},\bm{L},\bm{f},\bm{v},\bm{\psi},\bm{z}_{\lambda},\bm{z}_{\bm{\psi}})=\prod_{l=1}^{n}p(\lambda_{i}|\bm{\beta},\bm{\psi},z_{\lambda_{i}})$;
  6. 6.
    $p(\bm{\psi}|\bm{y},\bm{\beta},\bm{L},\bm{f},\bm{\Sigma},\bm{\lambda},\bm{z}_{\lambda},\bm{z}_{\bm{\psi}})=\prod_{i=1}^{n}\prod_{j=2}^{k}p(\psi_{i,j}|\beta_{i,j},\lambda_{i},z_{\psi_{i,j}})$;
  7. 7.
    $p(\bm{z}_{\lambda}|\bm{y},\bm{\beta},\bm{L},\bm{f},\bm{\Sigma},\bm{\lambda},\bm{\psi},\bm{z}_{\bm{\psi}})=\prod_{i=1}^{n}p(z_{\lambda_{i}}|\lambda_{i})$;
  8. 8.
    $p(\bm{z}_{\bm{\psi}}|\bm{y},\bm{\beta},\bm{L},\bm{f},\bm{\Sigma},\bm{\lambda},\bm{\psi},\bm{z}_{\lambda})=\prod_{i=1}^{n}\prod_{j=2}^{k}p(z_{\psi_{i,j}}|\psi_{i,j})$,
with $\bm{y}_{i}=(y_{i,1},\dots,y_{i,T})^{\prime}$ be a $T\times 1$ vector of observations of the $i-$th variable.
Step 1 First, we sample $\bm{f}_{t}$ for $t=1,\dots,T$. We can use standard regression results (see, e.g., Chan et al., (2019)) to obtain
$(\bm{f}_{t}|\bm{y},\bm{\beta},\bm{L})\sim N(\hat{\bm{f}}_{t},\bm{K}_{\bm{f}}^{-1}),$ (17)
where
$\bm{K}_{\bm{f}}^{-1}=(\bm{I}_{r}+\bm{L}^{\prime}\bm{\Sigma}\bm{L})^{-1},\quad\hat{\bm{f}}_{t}=\bm{K}_{\bm{f}}(\bm{L}\bm{\Sigma}^{-1}(\bm{y}_{t}-(\bm{I}_{n}\otimes\bm{x}^{\prime}_{t})\bm{\beta})).$ (18)
We use the efficient truncated Normal generator provided by Botev, (2017) in order to sample the factors.
Step 2 Second, we sample $\bm{L}$. Given the latent factors $\bm{f}$, the VAR becomes $n$ unrelated regressions and we can sample $\bm{L}$ equation by equation. Remember that $\bm{\beta}_{i}$ and $\bm{l}_{i}$ denote, respectively, the VAR coefficients and the factor loadings in the $i-$th equation. Then, the $i-$th equation of the VAR can be expressed as
$\bm{y}_{i}=\bm{X}_{i}\bm{\beta}_{i}+\bm{F}\bm{l}_{i}+\bm{v}$ (19)
where $\bm{F}=(\bm{f}_{1},\dots,\bm{f}_{r})$ the $T\times r$ matrix of factors with $\bm{f}_{i}=(f_{i,1},\dots,f_{i,T})^{\prime}$. The vector of noise $\bm{v}=(v_{i,1},\dots,v_{i,T})^{\prime}$ is distributed as $N(\bm{0},\bm{I}_{T}\sigma_{i}^{2})$.
Then using standard linear regression results, we get
$(\bm{l}_{i}|\bm{y}_{i},\bm{f},\bm{\beta}_{i},\sigma_{i}^{2})\sim N(\hat{\bm{l}_{i}},\bm{K}^{-1}_{\bm{l}_{i}})\bm{1}(\ell_{t}^{f}<\bm{R}_{t}^{f}\bm{f}_{t}<\upsilon_{t}^{f})\bm{1}(\ell_{it}^{Lf}<\bm{R}_{it}^{Lf}(\bm{l}_{i}^{\prime}\odot\bm{f}_{t})<\upsilon_{it}^{Lf},$ (20)
where
$\bm{K}_{\bm{l}_{i}}=\bm{V}^{-1}_{\bm{l}_{i}}+\bm{\sigma}^{-2}_{i}\bm{F}^{\prime}\bm{F},\quad\hat{\bm{l}}_{i}=\bm{K}^{-1}_{\bm{l}_{i}}(\bm{V}^{-1}_{\bm{l}_{i}}\bm{l}_{0,i}+\sigma^{-2}_{i}\bm{F}(\bm{y}_{i}-\bm{X}_{i}\bm{\beta}_{i})).$
We use the efficient truncated Normal generator provided by Botev, (2017) in order to sample $\bm{l}_{i}$.
Step 3 Third, we sample $\bm{\beta}$ equation by equation based on (19). Equation by equation estimation simplifies the estimation and allows for the estimation with a large number of variables. Again, using standard linear regression results, we get
$(\bm{\beta}_{i}|\bm{y}_{i},\bm{f},\bm{l}_{i},\sigma_{i}^{2})\sim N(\hat{\bm{l}_{i}},\bm{K}^{-1}_{\bm{\beta}_{i}}),$ (21)
where
$\bm{K}_{\bm{\beta}_{i}}=\bm{V}^{-1}_{\bm{\beta}_{i}}+\bm{\sigma}^{-2}_{i}\bm{X}_{i}^{\prime}\bm{X}_{i},\quad\hat{\bm{\beta}}_{i}=\bm{K}^{-1}_{\bm{\beta}_{i}}(\bm{V}^{-1}_{\bm{l}_{i}}\bm{\beta}_{0,i}+\sigma^{-2}_{i}\bm{X}_{i}(\bm{y}_{i}-\bm{F}\bm{l}_{i})).$
We use the efficient truncated Normal generator provided by Botev, (2017) in order to sample $\bm{l}_{i}$.
Step 4 Next, we sample $\sigma^{2}_{i}$ for $i=1,\dots,n$. Given $\bm{f}$ the model reduces to $n$ independent linear regressions. Therefore, we can use standard regression results (see, e.g., Chan et al., (2019)) to obtain
$(\sigma_{i}^{2}|\bm{y},\bm{f},\bm{L})\sim\mathcal{IG}\left(\alpha_{0}+\frac{T}{2},\beta_{0}+0.5\sum_{t=1}^{T}(y_{it}-\bm{X}_{it}\bm{\beta}_{i}-\bm{l}_{i}\bm{f}_{t})^{2}\right).$ (22)
Step 5 Lastly, we sample the hyperparameter $\lambda_{i}$ and $\psi_{i,j}$ from our shrinkage prior for the VAR coefficients as well as the mixing variables $z_{\lambda_{i}}$ and $z_{\psi_{i,j}}$. Using the latent variable representation of the half Cauchy distribution, we obtain
$p(\psi_{i,j}|\beta_{i,j},\lambda_{i},z_{\psi_{i,j}})$ $\propto\psi_{i,j}^{\frac{1}{2}}\text{e}^{-\frac{1}{2\lambda_{i}\psi_{i,j}}(\beta_{i,j}-m_{i,j})^{2}}\times\psi^{-\frac{3}{2}}\text{e}^{-\frac{1}{\psi_{i,j}z_{\psi_{i,j}}}}$
$=\psi_{i,j}^{-2}\text{e}^{-\frac{1}{\psi_{i,j}}\left(\frac{1}{z_{\psi_{i,j}}}+\frac{(\beta_{i,j}-m_{i,j})^{2}}{2\lambda_{i}}\right)},$
which is the kernel of the following inverse-gamma distribution:
$(\psi_{i,j}|\beta_{i,j},\lambda_{i},z_{\psi_{i,j}})\sim\mathcal{IG}\left(1,\frac{1}{z_{\psi_{i,j}}}+\frac{(\beta_{i,j}-m_{i,j})^{2}}{2\lambda_{i}}\right).$ (23)
Furthermore,
$p(\lambda_{i}|\bm{\beta},\bm{\psi},z_{\lambda_{i}})$ $\propto\prod_{i}\lambda_{i}^{-\frac{1}{2}}\text{e}^{-\frac{1}{2\lambda_{i}\psi_{i,j}}(\beta_{i,j}-m_{i,j})^{2}}\times\lambda_{i}^{-\frac{3}{2}}\text{e}^{-\frac{1}{\lambda_{i}z_{\lambda_{i}}}},$
$=\lambda_{i}^{-\left(\frac{k}{2}+1\right)}\text{e}^{-\frac{1}{\lambda_{i}}\left(\frac{1}{z_{\lambda_{i}}}+\sum_{j=2}^{k}\frac{(\beta_{i,j}-m_{i,j})^{2}}{2\psi_{i,j}}\right)},$
which is the kernel of the following inverse-gamma distribution:
$(\lambda_{i}|\bm{\beta},\bm{\psi},z_{\psi_{\lambda_{i}}})\sim\mathcal{IG}\left(\frac{k}{2},\frac{1}{z_{\lambda_{i}}}+\sum_{j=2}^{k}\frac{(\beta_{i,j}-m_{i,j})^{2}}{2\psi_{i,j}}\right).$ (24)
Finally, we sample the latent variables $\bm{z}_{\bm{\psi}}$ and $\bm{z}_{\bm{\lambda}}$. In particular, $z_{\psi_{i,j}}\sim\mathcal{IG}(1,1+\psi_{i,j}^{-1})$ for $i=1,\dots,n$ and $j=2,\dots,k$. Similarly, we have $z_{\lambda_{l}}\sim\mathcal{IG}(1,1+\lambda_{i}^{-1})$ for $i=1,\dots,n$.
3 Application to Stock and Watson, (2012)
This section presents an empirical application based on the framework of Stock and Watson, (2012), who estimate a 30-variable vector autoregression (VAR) to assess the relevance of five macroeconomic shocks during the Great Financial Crisis. This empirical application serves the purpose to illustrate the effectiveness of our proposed algorithm to combine impact and shock sign restrictions in large a structural VAR. In line with their approach, we estimate a large-scale VAR model comprising the same set of variables, simultaneously identifying five structural shocks: an oil supply (news) shock, a monetary policy shock, a technology shock, a financial risk shock, and a government spending shock. [^3] Our empirical strategy closely follows Hou, (2024), who also estimate the same 30-variable VAR and identify the aforementioned structural shocks using a proxy SVAR approach. Moreover, we label our oil shock as an oil supply shock. In contrast to Stock and Watson, (2012), we extend the sample period to span from 1983Q1 to 2019Q3 and employ both impact sign and shock sign restrictions for the identification of structural shocks. [^4] Breitenlechner et al., (2024) note that this sample period captures a relatively stable macroeconomic environment, excluding the heightened oil price volatility of the 1970s as well as the COVID-19 pandemic and the geopolitical shock resulting from the Russian invasion of Ukraine. A detailed overview of the 30 variables included in the VAR is provided in Section Online Appendix A. Following the transformations employed in Chan et al., (2025), most variables are expressed in logarithms. To mitigate the risk of overfitting in such a high-dimensional setting, we impose horseshoe priors, as proposed in Korobilis, (2022) and Chan et al., (2025) and described above. The baseline model is estimated with a lag length of $p=4$, which balances model complexity with parsimony. Posterior inference is conducted using the described Gibbs sampler. Specifically, we perform 6,000 iterations, discarding the first 1,000 as burn-in and retaining every 10th of the remaining draws for inference.
3.1 Impact Sign and Shock Sign Restrictions
In this section, we outline and discuss the implementation of impact sign and shock sign restrictions. Particular emphasis is placed on the shock sign restrictions for 2001Q3 and 2008Q4, given their role in disentangling structural shocks. The remaining shock sign restrictions are deferred to Appendix Online Appendix B.
Regarding the impact sign restrictions imposed on the $L$-matrix via $\ell_{i}^{L}<\bm{R}_{i}^{L}\bm{l}_{i}^{\prime}<\upsilon_{i}^{L}$ in Equation (7), we draw on a well-established literature on shock identification using sign restrictions. For the identification of a positive oil supply (news) shock, as presented in Table 1, we follow Banbura et al., (2023) and assume that such a shock leads to an increase in oil prices, as well as in both consumer and producer prices. [^5] Banbura et al., (2023) elaborate on the identification of oil supply and demand shocks, referring to Peersman and Van Robays, (2009), Kilian, (2008), Kilian, (2009), Kilian and Murphy, (2014), and Baumeister and Hamilton, (2019). Notably, our oil supply news is closely related to the shock identified by of Oil Supply News: Evidence from OPEC Announcements, (2021), where a positive shock implies the expectation of lower oil supply. In addition, we assume that a positive oil supply shock induces a monetary tightening, depresses stock prices, and worsens labor market conditions. Output and its components are left unrestricted. Concerning monetary policy shocks, we adopt the identification strategy of Chan et al., (2025), under which a tightening of monetary policy is associated with increases in interest rates and bond yields, declines in stock prices, and a contraction in economic activity. [^6] Extensive discussions on the identification of monetary policy shocks in VAR frameworks can be found in Sims, (1972), Sims, (1980), Sims, (1986), Faust, (1998), Canova and Pina, (2005), Canova and De Nicolo, (2002), Uhlig, (2005), and Rubio-Ramirez et al., (2010). The technology shock is identified following Chan et al., (2025), with the assumption that it increases economic activity, raises labor demand, and exerts downward pressure on prices. In response to such a shock, stock markets are expected to appreciate, while the policy rate declines in reaction to the deflationary environment. [^7] For additional discussion on the sign identification of technology shocks, see Dedola and Neri, (2007) and Peersman and Straub, (2009). To identify a financial risk shock, we assume a widening of the yield spread between corporate and ten year government bonds, a decline in stock market valuations, and a decrease in consumer prices, an approach consistent with Chan et al., (2025). [^8] See also Korobilis, (2022) and Furlanetto et al., (2019) for further discussion on the identification of financial shocks. Finally, a government spending shock is identified by assuming an increase in government expenditures, a reduction in unemployment, and a rise in (stock) prices. This identification strategy also aligns with Chan et al., (2025). [^9] See Laumer, (2020) and Forni et al., (2010) for detailed treatments of fiscal shock identification.
As previously noted, we impose shock sign restrictions on the oil supply, financial risk, and government spending shocks in 2001Q3 and 2008Q4 through $\ell_{\tilde{t}}^{f}<\bm{R}_{\tilde{t}}^{f}\bm{f}_{\tilde{t}}<\upsilon_{\tilde{t}}^{f}$ in Equation (8), as these restrictions are essential for the separation of the respective shocks. In 2001Q3, the terrorist attacks on September 11th resulted in a sudden and pronounced increase in financial risk, reflecting heightened uncertainty across financial markets. The temporary closure of the New York Stock Exchange and the flight of investors toward safe-haven assets such as Treasury bonds substantiate this assessment. [^10] This interpretation is further supported by the Financial Stress Index; see Federal Reserve Bank of St. Louis, (2025). In terms of the oil shock, we assume it to be positive, as uncertainty surrounding oil supply rose significantly following the attacks, with market participants anticipating potential production cuts and hence, higher oil prices. Additionally, we assume a positive government spending shock in response to the enactment of the Emergency Supplemental Appropriations Act on September 18th, 2001.111111For further details, see Congress, (2001), Ramey and Zubairy, (2018), and Blanchard and Perotti, (2002). For the fourth quarter of 2008, we impose a negative oil supply news shock, a positive financial risk shock, and a negative government spending shock. The bankruptcy of Lehman Brothers in late September 2008 precipitated widespread turmoil in global financial markets, leading to a significant increase in financial uncertainty during the subsequent quarter. Consequently, we posit a positive financial risk shock.  [^12] See again Federal Reserve Bank of St. Louis, (2025) for empirical evidence. Regarding the oil supply shock, the financial crisis contributed to a downward revision of oil price expectations, thereby alleviating supply constraints.  [^13] In fact, the Organization of the Petroleum Exporting Countries (OPEC) held two meetings in 2008Q4 to address the decline in oil prices (Organization of the Petroleum Exporting Countries, , 2025). In terms of government spending, the election of President Obama resulted in a reduced expectation of government expenditure due to announcements of cuts in military spending.  [^14] Additionally, the American Recovery and Reinvestment Act, aimed at stimulating the economy, was announced by President Obama in January 2009 and signed into law in February 2009 (Congress, , 2009). As illustrated in Table 1, the narrative shock sign restrictions for both the financial risk and government spending shocks are critical for disentangling the two shocks. The additional shock sign restrictions further refine the identification process. It is noted that the monetary policy and the financial risk shock are not separated by means of the impact sign or shock sign restrictions. However, and as seen below in 3.2 and Online Appendix C the shock sign restrictions suffice. Further, we check the uniqueness of the sign patterns of the every retained shock impact matrix $L$ and the structural shocks $f_{t}$. [^15] The sign restrictions on shocks over distinct time periods may be interpreted as a more agnostic framework than impact sign restrictions, as noted in Antolín-Díaz and Rubio-Ramírez, (2018). Nonetheless, it is emphasized that sign restrictions must be grounded in a plausible narrative foundation.
Table 1: Impact Sign and Shock Sign Restrictions to Identify Structural Shocks
Impact Sign Restrictions Variable Oil Supply Monetary Policy Technology Financial Risk Government Spending GDP $-1$ $+1$ Consumption $-1$ $+1$ Investment $-1$ Gov. Spending $+1$ Hours $-1$ Real Compensation $+1$ $-1$ Output per Hour $+1$ Unit Labor Cost $+1$ $+1$ Industrial Production $-1$ Capacity Utilization Employment $-1$ $-1$ $+1$ Unemployment $+1$ $+1$ $-1$ $-1$ Housing Starts M&T Sales M&T Inventories PCE Index $+1$ $-1$ $-1$ $-1$ $+1$ GDP Deflator $+1$ $-1$ $-1$ $-1$ $+1$ CPI $+1$ $-1$ $-1$ $-1$ $+1$ PPI $+1$ $-1$ FED Funds Rates $+1$ $+1$ $-1$ 3-Month TB $+1$ 1-Year TB $+1$ 10-Year TB $+1$ BAA-GS10 Spread $-1$ $+1$ Monetary Base $-1$ M2 $-1$ S&P500 $-1$ $-1$ $+1$ $-1$ $+1$ Dow Jones $-1$ $-1$ $+1$ $-1$ $+1$ US Dollar Index $+1$ Oil Price $+1$ $+1$ Narrative Shock Sign Restrictions Oil Supply Monetary Policy Technology Financial Risk Government Spending $>0$ 1986Q3, 1988Q4 2001Q3, 2016Q4 2001Q3, 2008Q4 1990Q4, 2001Q3, 2001Q4, 2002Q1, 2002Q3, 2003Q1, 2006Q2, 2007Q4 $<0$ 2001Q4, 2008Q4 2014Q4 1998Q4, 1999Q4 1986Q3, 1988Q4, 1989Q4, 1991Q4, 2008Q4, 2011Q3, 2013Q1 Notes: This table summarizes the impact sign restrictions and narrative shock sign restrictions to identify the five structural shocks. In the upper table, $+1$ ($-1$) implies a positive (negative) on impact response of the respective variable in the left column. In the lower table, $>0$ ($<0$) restricts the respective shock in the column to be positive (negative).
3.2 Empirical Results
In this section, we present selected results from our empirical analysis. Figure 1 displays the impulse response functions (IRFs) of key macroeconomic and financial variables, including real GDP, industrial production, unemployment, the Consumer Price Index (CPI), and the S&P 500 index, in response to identified structural shocks. The oil supply news shock, which manifests as an increase in oil prices, leads to a contraction in economic activity, a decline in labor demand, and an increase in consumer prices. As expectations about future economic conditions deteriorate, equity markets also decline. These dynamics are consistent with the empirical findings in the literature (e.g., of Oil Supply News: Evidence from OPEC Announcements, (2021), Kilian and Park, (2009), Kilian, (2009), Hamilton, (2003), Ramey and Vine, (2011)). A contractionary monetary policy shock is found to reduce output, raise unemployment, and lower both prices and equity valuations (Christiano et al., (2005), Gertler and Karadi, (2015), Jarocinski and Karadi, (2020)). This response pattern suggests that the identified monetary policy shock is purged from any information effects or central bank signaling, thereby capturing a pure policy shock (see Jarocinski, (2022), Bauer and Swanson, 2023a , Bauer and Swanson, 2023b ). A positive technology shock increases both output and labor demand, while exerting downward pressure on prices due to productivity improvements. The stock market responds favorably. This is largely in line with existing evidence on the expansionary effects of technological innovations (Peersman and Straub, (2009)). A shock to financial risk has strongly contractionary effects: output and prices decline persistently, unemployment rises, and stock markets fall. These results corroborate prior findings on the macroeconomic consequences of financial stress (Gilchrist and ZakrajÅ¡ek, (2012), Caldara et al., (2016), Christiano et al., (2014), Jurado et al., (2015), Korobilis, (2022)). Finally, an increase in government spending yields an expansion in output, a reduction in unemployment, upward pressure on prices, and a short-run increase in equity prices. These responses are consistent with a substantial body of empirical work (Blanchard and Perotti, (2002), Ramey and Shapiro, (1998), Ramey, (2011), Auerbach and Gorodnichenko, (2012), Mountford and Uhlig, (2009), Laumer, (2020)). The complete set of impulse responses is provided in Section Online Appendix C of the Online Appendix.
Figure 1: Selected Impulse Responses
Notes: Median impulse responses of selected variables to an oil supply, monetary policy, technology, financial risk and government spending shock. The solid red line depict the medians and the shaded red areas the 68% credible bands.
Table 2 presents the forecast error variance decomposition (FEVD) of GDP across various forecast horizons. The FEVD illustrates the proportion of the variance in GDP forecast errors that can be attributed to each of the identified structural shocks. Across all horizons, financial risk shocks emerge as the dominant driver of GDP fluctuations. Compared to the findings of Hou, (2024), our results suggest a greater role for oil supply shocks in driving the business cycle. This divergence may stem from the extended sample period employed in our empirical analysis. [^16] It should be noted that due to differences in identification strategies, our results are not directly comparable to those of Hou, (2024). Moreover, their sample is restricted to the period from 1987 to 2011, which is heavily influenced by the global financial crisis. In contrast, government spending shocks exhibit only a moderate contribution to the forecast error variance of GDP, while monetary policy and technology shocks account for a relatively smaller share of GDP fluctuations across all horizons. In the Online Appendix, Section Online Appendix C shows the historical decomposition of GDP corroborating the finding that financial risk shocks are the largest driver of the business cycle. Further, in Section Online Appendix C shows that if we shorten our sample and end in 2012 financial risk shocks take on an even more important role in driving the business cycle. [^17] Further results on the above-explained shortened sample can be obtained on request.
Table 2: Forecast error variance decompositions of GDP
Oil Supply Monetary Policy Technology Financial Risk Government Spending
$H=1$ 28% 2% 4% 57% 8%
$H=5$ 28% 2% 5% 57% 8%
$H=10$ 29% 3% 6% 55% 9%
$H=20$ 29%1 3% 7% 53% 9%
Notes: Table presents posterior medians of the forecast error variance decompositions of GDP attributed to the oil supply, monetary policy, technology, financial risk, and government spending shocks at horizons $H=1$, $H=5$, $H=10$ and $H=20$.
4 Conclusion
This paper presents a novel high-dimensional SVAR framework that allows for the imposition of a large number of linear inequality restrictions on both impact impulse responses and structural shocks, including their interactions. By leveraging a factor structure and introducing a computationally efficient sampling algorithm that avoids rejection-based procedures, our approach overcomes key limitations of existing methods, most notably their poor scalability in high dimensional settings. The flexibility to combine sign and shock restrictions enhances the precision of structural shock identification and allows for economically meaningful inference even in large systems.
Our empirical application, identifying five structural shocks in a 30 variable VAR, demonstrates the practical relevance of the methodology, particularly in uncovering the prominent role of financial shocks in driving business cycle fluctuations. As regards the narrative shock sign restrictions, we illustrate their capacity to sharpen inference by shrinking the identification set. More importantly, shock sign restrictions can facilitate the disentangling of structural shocks as shown for the financial risk and government risk shock. Overall, the proposed framework significantly broadens the scope for credible and computationally feasible structural analysis in large VAR environments.
Appendix Online Appendix A Data Set - Empirical Application
Table A.1: List of the 30 Macroeconomic Variables Used in the Empirical Application
Variable Description FRED CODE
GDP Real Gross Domestic Product GDPC1
Consumption Real Personal Consumption Expenditures PCECC96
Investment Real Gross Private Domestic Investment GPDIC1
GovSpending Real Government Consumption Expenditures and Gross Investment GCEC1
EmpHur Nonfarm Business Sector: Hours of All Persons HOANBS
CPH Nonfarm Business Sector: Real Compensation Per Hour COMPRNFB
OPH Business Sector: Output Per Hour of All Persons OPHNFB
ULC Business Sector: Unit Labor Cost ULCBS
IP Industrial Production: Total index INDPRO
CapuTot Capacity Utilization: Total Industry TCU
Employment Total Nonfarm Payrolls: All Employees PAYEMS
Unemployment Unemployment Rate UNRATE
HStarts Housing Starts: Total: New Privately Owned Housing Units Started HOUST
MTSales Manufacturing and trade sales CMRMTSPL
MTInvent Manufacturing and trade inventories INVCMRMTSPL
PCED Personal Consumption Expenditures: Chain-type Price Index PCEPI
GDPDefl Gross Domestic Product: Chain type Price Index GDPCTPI
CPI Consumer Price Index For All Urban Consumers: All Items CPIAUCSL
PPI Producer Price Index: All Commodities PPIACO
FedFunds Eective Federal Funds Rate FEDFUNDS
TB3Mth 3-Month Treasury Bill: Secondary Market Rate TB3MS
TB1Yr 1-Year Treasury Constant Maturity Rate GS1
TB10Yr 10-Year Treasury Constant Maturity Rate GS10
BAAGS10Spread BAA-GS10 spread BAA - GS10
MBase Monetary Base Adjusted for changes in Reserve Requirement BOGMBASE
M2 M2SL M2SL
SP500 S&P’s Common Stock Price Index: Composite
DJIA Common Stock Prices: Dow Jones Industrial Average
DollarIndex FRB Nominal Major Currencies Dollar Index TWEXMMTH & TWEXAFEGSMTH
Oil Price PPI: Crude Petroleum WPU0561
Notes: This table lists the 30 macroeconomic variables used in the empirical application, along with their descriptions and FRED source codes. SP500 and DJIA are retrieved from LSEG Data & Analytics. All variables, except the federal funds rate, 3-Month treasury bill rate, 1-year treasury bill rate, 10-year treasury bill rate and the corporate bond spread are transformed to logs.
Appendix Online Appendix B Narrative Shock Sign Restrictions
In this Appendix Section, we will outline the narrative shock sign restrictions for the oil supply, financial risk and government spending shocks that we did not discuss in the Section 3.1.
Oil Supply
1986Q3: In August 1986, OPEC members reached a consensus to reinstate production quotas, aiming to reduce the cartel’s output by nearly 4 million barrels per day. The announcement of this agreement had an immediate impact on market expectations as it increased oil price expectation. We assume the shock to be positive.
1988Q4: In late November 1988, OPEC members reached a unanimous agreement to reduce crude oil production, setting a ceiling of 18.5 million barrels per day effective January 1, 1989. This decision aimed to address the oversupply in the market and stabilize falling oil prices. We assume the shock to be positive.
2001Q4: On November 14, 2001, amid a global economic slowdown intensified by the September 11 attacks, OPEC announced it would only cut oil production if non-OPEC producers did the same. This conditional stance was interpreted by markets as a potential signal of failed coordination and the risk of a price war, prompting a sharp downward revision in oil price expectations (Al-Naimi, (2016), of Oil Supply News: Evidence from OPEC Announcements, (2021)). We assume the shock to be negative.
2014Q4: A significant downward revision of oil price expectations occurred on November 27, 2014, when OPEC announced it would maintain existing production levels. Prior to the meeting, many market participants had anticipated that the cartel would implement production cuts to support falling oil prices. However, Saudi Arabia opposed calls from several economically weaker OPEC members to reduce output, resulting in a roughly 10 percent drop in price expectations (of Oil Supply News: Evidence from OPEC Announcements, , 2021). We assume the shock to be negative.
2016Q4: On November 30, 2016, OPEC announced its first oil production cut since 2008, agreeing to reduce output by 1.2 million barrels per day, setting a new production ceiling of 32.5 million barrels per day effective January 1, 2017. This decision was aimed at addressing the global supply glut and stabilizing falling oil prices. We assume the shock to be positive.
Financial Risk
1998Q4: In the fourth quarter of 1998, global financial markets showed clear signs of stabilization after enduring months of heightened volatility and turmoil triggered by the Asian Financial Crisis. The crisis, which began in 1997, had a profound impact on emerging market economies, particularly in Southeast Asia, and spread to other parts of the world, triggering capital flight, currency devaluations, and a series of banking crises. We assume the shock to be negative.
1999Q4: In Q4 1999, global financial risk decreased as markets continued to recover from the Asian Financial Crisis and the Russian default in 1998. Investor confidence in emerging markets, particularly in Southeast Asia, was bolstered by successful reforms and international support. The U.S. economy showed strong growth, supported by low interest rates and a stable financial sector, further reducing systemic risk. Additionally, global stock markets, particularly in the U.S. and Europe, experienced steady gains, while the dot-com boom contributed to an overall optimistic outlook, leading to a decrease in perceived financial risks. We assume the shock to be negative.
Government Spending
1986Q3: This quarter is best understood as reflecting a negative narrative shock due to news of tighter future fiscal constraints and a softening in Cold War urgency, which together implied lower future military spending. We assume the shock to be negative.
1988Q4: The quarter saw the election of George H. W. Bush (November 1988), who, while aligned with Reagan’s Republican platform, signaled greater fiscal restraint and pragmatism. There was growing bipartisan concern over persistent federal deficits, setting the stage for tighter budget policies in the future, including scrutiny of defense spending. We assume the shock to be negative.
1989Q4: The negative government spending shock in 1989Q4 can be attributed to a sudden and largely unanticipated geopolitical transformation, most notably the collapse of the Eastern Bloc and the symbolic end of the Cold War. These developments led to a significant downward revision in expectations for future U.S. defense expenditures (peace dividend).
1990Q4: The positive spending shock in 1990Q4 reflects the unexpected shift from post-Cold War military contraction to renewed engagement, driven by the rapid mobilization and political commitment surrounding the Gulf crisis. This period marked a clear reversal of previous expectations, with news pointing toward significantly higher future military spending.
1991Q4: The negative government spending shock reflects the post-conflict drawdown, the collapse of the Soviet Union, and a renewed policy focus on deficit reduction and demilitarization. These factors generated news of lower expected future military spending, reinforcing the shift toward a post-Cold War fiscal and security environment.
2001Q4: The positive government spending shock in was driven by the unexpected and large-scale revision in defense and security spending expectations following the 9/11 attacks and the launch of the War on Terror. The perceived need for a sustained military response, global engagement, and enhanced domestic security produced a sharp upward shift in anticipated future government expenditure.
2002Q1: In early 2002, the U.S. experienced a positive government spending shock driven by rising expectations of expanded military engagement following 9/11. President Bush’s “Axis of Evil” speech in January signaled potential future conflicts beyond Afghanistan, particularly with Iraq. This shift in rhetoric increased the perceived likelihood of sustained military operations. At the same time, the administration proposed the largest defense budget increase since the 1980s in its FY2003 submission, reflecting a structural rise in defense spending. Together, these developments contributed to a sharp upward revision in expected government military outlays.
2002Q3: In 2002Q3, the U.S. experienced a positive government spending shock driven by growing momentum toward military intervention in Iraq. Throughout the quarter, the Bush administration intensified its public case for action, culminating in a push for congressional authorization to use force, which signaled an increased likelihood of large-scale military deployment. In August and September, leaked reports and official statements suggested that war planning was well underway, with expectations of a buildup in troops and materiel. These developments represented new information about future military operations and contributed to a further upward revision in anticipated defense spending.
2003Q1: The positive government spending shock in 2003Q1 was driven by credible and increasingly certain news of an impending war in Iraq, alongside budget signals pointing to a sustained rise in defense expenditures. The scale of troop deployments, political messaging, and preemptive budget actions all contributed to a marked upward revision in expected future military spending.
2003Q4: The positive government spending shock in 2003Q4 reflects growing awareness that the costs of the Iraq War would be significantly higher and more persistent than previously expected. The \$87 billion supplemental package and the worsening security situation contributed to a sharp upward revision in anticipated future military spending.
2004Q2: In 2004Q2, U.S. government spending rose significantly, primarily due to the ongoing military operations in Iraq and Afghanistan. The Bush administration had requested a large supplemental budget to support these operations, which included funding for troop maintenance, logistics, and equipment.
2005Q1: In 2005Q1, U.S. government spending increased again due to the military operations in Iraq and Afghanistan. A key event during this period was President Bush’s request for an emergency supplemental budget, which included funding to maintain U.S. forces in Iraq and Afghanistan as well as support for reconstruction efforts in Iraq.
2006Q2: In 2006Q2, U.S. government spending rose notably, largely due to the continued military operations in Iraq and Afghanistan. A key event was President Bush’s request for additional funds through a supplemental budget, aimed at supporting military efforts and rebuilding Iraq.
2007Q4: In 2007Q4, U.S. government spending saw a notable increase, largely driven by continued military engagement in Iraq and Afghanistan. During this period, the U.S. government authorized substantial funding to support the ”surge” strategy in Iraq, which involved deploying additional troops and resources to stabilize the country.
2011Q3: In 2011Q3, U.S. government spending experienced a decline, which can be attributed to a shift in military expenditures following the drawdown of U.S. forces in Iraq. A significant event was the announcement by President Obama of a planned withdrawal of combat troops from Iraq, signaling a reduction in military spending associated with the war.
2013Q1: In 2013Q1, U.S. government spending decreased, largely driven by the implementation of the Budget Control Act of 2011, which triggered automatic spending cuts, known as sequestration. The sequestration cuts affected military and defense spending, leading to a reduction in funds allocated for ongoing operations, including those in Afghanistan.
Appendix Online Appendix C Additional Empirical Results
Impulse Responses
Figure OC.1: Oil Supply News Shock
Notes: Median impulse responses to an oil supply shock. The solid red lines depict the medians and the shaded red areas the 68% credible bands.
Figure OC.2: Monetary Policy Shock
Notes: Median impulse responses to a monetary policy shock. The solid red lines depict the medians and the shaded red areas the 68% credible bands.
Figure OC.3: Technology Shock
Notes: Median impulse responses to a technology shock. The solid red lines depict the medians and the shaded red areas the 68% credible bands.
Figure OC.4: Financial Risk Shock
Notes: Median impulse responses to a financial risk shock. The solid red lines depict the medians and the shaded red areas the 68% credible bands.
Figure OC.5: Government Spending Shock
Notes: Median impulse responses to a government spending shock. The solid red lines depicts the medians and the shaded red areas the 68% credible bands.
Historical Decomposition of GDP
Notes: Posterior medians of the historical contributions of the oil supply, monetary policy, technology, financial risk, and government spending shocks to GDP.
Sample Period 1983 to 2011
Figure OC.6: Selected Impulse Responses
Notes: Median impulse responses of selected variables to an oil supply, monetary policy, technology, financial risk and government spending shock. The solid red line depict the medians and the shaded red areas the 68% credible bands (Sample period: 1983 to 2011).
Table C.2: Forecast error variance decompositions of GDP
Oil Supply Monetary Policy Technology Financial Risk Government Spending
$H=1$ 5% 13% 5% 68% 9%
$H=5$ 5% 14% 5% 66% 9%
$H=10$ 6% 17% 5% 62% 10%
$H=20$ 9%1 17% 7% 55% 12%
Notes: Table presents posterior medians of the forecast error variance decompositions of GDP attributed to the oil supply, monetary policy, technology, financial risk, and government spending shocks at horizons $H=1$, $H=5$, $H=10$ and $H=20$ (Sample period: 1983 to 2011).